OpenStudy (anonymous):

rewrite the expression so it is not in fractional form sin^2 y/1 − cos y

OpenStudy (phi):

try multiplying top and bottom by (1+cos y)

OpenStudy (anonymous):

I still can't figure this problem out

OpenStudy (anonymous):

\[\frac{ \sin^2(y) }{ 1-\cos(y) }\]

OpenStudy (phi):

what do you get if you multiply the bottom by 1 + cos y ? (and to keep things kosher, also multiply the top by 1 + cos y )

OpenStudy (phi):

notice (a-b)(a+b) = a^2 - b^2 is the "short" way to do the multiply

OpenStudy (anonymous):

do you get 1 if you multiply them?

OpenStudy (phi):

no. do you know FOIL ? that is the long way (1 - cos y ) ( 1 + cos y) First outer inner last

OpenStudy (anonymous):

oh wow why didn't that come to me the first 50 times I tried this problem?

OpenStudy (anonymous):

so then after foiling would I be left with 1-cos^2(y)?

OpenStudy (phi):

yes, and there is this *very useful* identity sin^2 y + cos^2 y = 1 from which we get sin^2 y = 1 - cos^2 y

OpenStudy (anonymous):

Thank you so much!

OpenStudy (phi):

you should get \[ \frac{\sin^2 y\ (1+ \cos y)}{\sin^2 y} = 1+\cos y\]

OpenStudy (anonymous):

That's exactly what I got, thank you for your help!